# On the Curses of Future and History in Future-dependent Value Functions for Off-policy Evaluation

### v1

@@ 5.3 New Algorithm
The discovery of effective history weights and its boundedness also leads to a new algorithm analogous to MIS methods for MDPs (Uehara et al., 2020). The idea is that since Lemma 1 tells us to minimize $\mathbb{E}_{\pi_{e}}[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})]$, which equals $\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})]$ from Eq.(10), we can then use another function class $\mathcal{W}\subset(\mathcal{H}\to\mathbb{R})$ to model $w^{\star}$, and approximately solve the following:
@@
$\inf_{w\in\mathrm{sp}(\mathcal{W})}\max_{V\in\mathcal{V}}\left|\sum_{h=1}^{H}\mathbb{E}_{\pi_{b}}\left[(w_{h}^{\star}(\tau_{h})-w_{h}(\tau_{h}))\cdot(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})\right]\right|$
$\qquad\leq\epsilon_{\mathcal{W}}.$
@@
We present the sample complexity bound of our algorithm as follows. Similar to Theorem 9, we assume that $C_{\mathcal{V}}:=\max_{V\in\mathcal{V}}\|V\|_{\infty}$ and $C_{\mathcal{W}}:=\max_{h}\sup_{w\in\mathcal{W}}\|w\|_{\infty}$ are not much larger than the corresponding norms of $V_{\mathcal{F}}$ and $w^{\star}$, respectively. See the proof in Appendix D.10.
@@
Let $\widehat{V}$ be the result of approximating Eq.13 with empirical estimation in Eq.(14). Assume that $C_{\mathcal{V}}\leq c\|V_{\mathcal{F}}\|_{\infty}$ for $V_{\mathcal{F}}$ in Eq.(12), and $C_{\mathcal{W}}\leq c\|w^{\star}\|_{\infty}$ for $w^{\star}$ in Eq.(11). Under Assumptions 11, 12, and 13, w.p. $\geq 1-\delta$, $\left|J(\pi_{e})-\mathbb{E}_{\mathcal{D}}[\widehat{V}(f_{1})]\right|\leq$
$\epsilon_{\mathcal{V}}+\epsilon_{\mathcal{W}}+cH^{2}C_{\mathcal{H},\infty}(C_{\mathcal{F},\infty}+1)\sqrt{\frac{C_{\mu}\log\frac{|\mathcal{V}||\mathcal{W}|}{\delta}}{n}}.$
@@ 6 Conclusion and Future Work
We conclude the paper with open questions for future work.
Curse of Memory  Our work considers memoryless policies. Similar to Uehara et al. (2022), it is possible to extend to policies that depend on recent observations and actions (or memory) $z_{h}$ by replacing latent $s_{h}$ with $(s_{h},z_{h})$ in various places (e.g., rows of $M_{\mathcal{H},h}$ and $M_{\mathcal{F},h}$). However, the size of $(s_{h},z_{h})$ grows exponentially with the length of $z_{h}$, which we call the curse of memory. We suspect that structural assumptions on the policies are needed to avoid the curse of memory, and leave this to future work.
@@ Acknowledgements
NJ thanks Yu Bai, Akshay Krishnamurthy, and Masatoshi Uehara for discussions on an early draft of the work.
Appendix A Discussion
A.1 Connection between Future-dependent Value Functions and Learnable Future-dependent Value Functions
Uehara et al. (2022) proposed a learnable version of future dependent value functions and focused on the learning of learnable FDVFs instead of the original FDVFs defined in Definition 3. In this subsection, we first introduce the definition of learnable FDVFs and then show that it is equivalent to FDVFs under a full-rank assumption.
@@ A.1 Connection between Future-dependent Value Functions and Learnable Future-dependent Value Functions
A.2 Comparison between the Finite-horizon and the Discounted Formulations
The original results of Uehara et al. (2022) were given in the infinite-horizon discounted setting. To describe their data collection assumption in simple terms, consider an infinite trajectory that enters the stationary configuration of $\pi_{b}$, and we pick a random time step and call it $t=0$ (which corresponds to a time step $h$ in our setting). Then, history (analogous to our $\tau_{h}$) is $(o_{(-M_{F}):(-1)},a_{(-M_{F}):(-1)})$, and future is $(o_{0:M_{H}},a_{0:M_{F}-1})$. $M_{F}$ and $M_{H}$ are hyperparameters that determines the lengths, and should not be confused with our $M_{\mathcal{F},h}$ and $M_{\mathcal{H},h}$ matrices. Data points collected in this way form the “transition” dataset. Separately, there is an “initial” dataset of $(o_{0:M_{F}-1},a_{0:M_{F}})$ tuples to represent the analogue of initial state distribution in MDPs.
@@ A.3 Refined Coverage
Appendix B Proofs for Section 3
B.1 Proof of Lemma 1

### v2

@@ 5.3 New Algorithm
The discovery of effective history weights and its boundedness also lead to a new algorithm analogous to MIS methods for MDPs (Uehara et al., 2020). The idea is that since Lemma 1 tells us to minimize $\mathbb{E}_{\pi_{e}}[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})]$, which equals $\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})]$ from Def 10, we can then use another function class $\mathcal{W}\subset(\mathcal{H}\to\mathbb{R})$ to model $w^{\star}$, and approximately solve the following:
@@
$\inf_{w\in\mathrm{sp}(\mathcal{W})}\max_{V\in\mathcal{V}}\left|\sum_{h=1}^{H}\mathbb{E}_{\pi_{b}}\left[(w_{h}^{\star}(\tau_{h})-w_{h}(\tau_{h}))\cdot(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})\right]\right|\qquad\leq\epsilon_{\mathcal{W}}.$
@@
We present the sample complexity bound of our algorithm as follows. Similar to Theorem 7, we assume that $C_{\mathcal{V}}:=\max_{V\in\mathcal{V}}\|V\|_{\infty}$ and $C_{\mathcal{W}}:=\max_{h}\sup_{w\in\mathcal{W}}\|w\|_{\infty}$ are not much larger than the corresponding norms of $V_{\mathcal{F}}$ and $w^{\star}$, respectively. See the proof in Appendix E.11.
@@
Let $\widehat{V}$ be the result of approximating Eq.9 with empirical estimation in Eq.(10). Assume that $C_{\mathcal{V}}\leq c\|V_{\mathcal{F}}\|_{\infty}$ for $V_{\mathcal{F}}$ in Eq.(5), and $C_{\mathcal{W}}\leq c\|w^{\star}\|_{\infty}$ for $w^{\star}$ in Eq.(8). Under Assumptions 12, 9, and 13, w.p. $\geq 1-\delta$, $\left|J(\pi_{e})-\mathbb{E}_{\mathcal{D}}[\widehat{V}(f_{1})]\right|\leq\epsilon_{\mathcal{V}}+\epsilon_{\mathcal{W}}+cH^{2}C_{\mathcal{H},\infty}(C_{\mathcal{F},\infty}+1)\sqrt{\frac{C_{\mu}\log\frac{|\mathcal{V}||\mathcal{W}|}{\delta}}{n}}.$
@@ 6 Conclusion and Future Work
The main text considers memoryless policies. Similar to Uehara et al. (2022a), we can extend to policies that depend on recent observations and actions (or memory). In fact, we provide a more general result in Appendix B.6 that handles recurrent policies that are finite state machines, which allows the policy to depend on long histories. However, the coverage coefficient will be diluted quickly when the memory contains rich information, which we call the curse of memory. We suspect that structural policies are needed to avoid the curse of memory and leave this to future work.
@@ Acknowledgements
Nan Jiang acknowledges funding support from NSF IIS-2112471, NSF CAREER IIS-2141781, Google Scholar Award, and Sloan Fellowship.
Appendix A Related Literature
OPE in POMDPs.
There is a line of research on OPE in confounded POMDPs [Zhang and Bareinboim, 2016, Namkoong et al., 2020, Nair and Jiang, 2021, Guo et al., 2022, Lu et al., 2022, Shi et al., 2022, Hong et al., 2023]. However, all these methods require the behavior policy to solely depend on the latent state. This assumption is inapplicable for our unconfounded POMDP setting, where the behavior policy depends on the observations. Hu and Wager [2023] studied the same unconfounded setting as ours. Nonetheless, their method uses multi-step importance sampling, leading to an undesirable exponential dependence on the horizon. The closest related work to our paper is [Uehara et al., 2022a], which we analyze in detail in Section 3.
Online Learning in POMDPs.
There are many prior works studying online learning algorithms for sub-classes of POMDPs, including decodable POMDPs [Krishnamurthy et al., 2016, Jiang et al., 2017, Du et al., 2019, Efroni et al., 2022], Latent MDPs [Kwon et al., 2021] and linear quadratic Gaussian setting (LQG) [Lale et al., 2021, Simchowitz et al., 2020]. Recently, online learning algorithms with polynomial sample results have been proposed for both tabular POMDPs [Guo et al., 2016, Azizzadenesheli et al., 2016, Jin et al., 2020, Liu et al., 2022a] and POMDPs with general function approximation [Zhan et al., 2022, Liu et al., 2022b, Chen et al., 2022, Huang et al., 2023]. All of these approaches are model-based and require certain model assumptions. Uehara et al. [2022b] proposed a model-free online learning algorithm based on future-dependent value functions (FDVFs). However, their algorithm requires the boundness of FDVFs for all policies in the policy class, as well as a low-rank property of the Bellman loss, which limits the generality of the results.
OPE in MDPs.
There has been a long history of studying OPE in MDPs, including importance sampling (IS) approaches [Precup et al., 2000, Li et al., 2011] and their doubly robust variants [Dudík et al., 2011, Jiang and Li, 2016, Thomas and Brunskill, 2016, Farajtabar et al., 2018], marginalized importance sampling (MIS) methods [Liu et al., 2018, Xie et al., 2019, Kallus and Uehara, 2020] and model-based estimators [Eysenbach et al., 2020, Yin et al., 2021, Voloshin et al., 2021]. As mentioned in Section 1, directly applying IS-based approaches for POMDPs will result in exponential variance. Meanwhile, MIS-based approaches do not apply to POMDPs since they require the environment to satisfy the Markov assumption.
Appendix B Discussions and Extensions
B.1 Connection between Future-dependent Value Functions and Learnable Future-dependent Value Functions
Uehara et al. [2022a] proposed a learnable version of future dependent value functions and focused on the learning of learnable FDVFs instead of the original FDVFs defined in Definition 3. In this subsection, we first introduce the definition of learnable FDVFs and then show that it is equivalent to FDVFs under a full-rank assumption.
@@ B.1 Connection between Future-dependent Value Functions and Learnable Future-dependent Value Functions
B.2 Comparison between the Finite-horizon and the Discounted Formulations
The original results of Uehara et al. [2022a] were given in the infinite-horizon discounted setting. To describe their data collection assumption in simple terms, consider an infinite trajectory that enters the stationary configuration of $\pi_{b}$, and we pick a random time step and call it $t=0$ (which corresponds to a time step $h$ in our setting). Then, history (analogous to our $\tau_{h}$) is $(o_{(-M_{F}):(-1)},a_{(-M_{F}):(-1)})$, and future is $(o_{0:M_{H}},a_{0:M_{F}-1})$. $M_{F}$ and $M_{H}$ are hyperparameters that determines the lengths, and should not be confused with our $M_{\mathcal{F},h}$ and $M_{\mathcal{H},h}$ matrices. Data points collected in this way form the “transition” dataset. Separately, there is an “initial” dataset of $(o_{0:M_{F}-1},a_{0:M_{F}})$ tuples to represent the analogue of initial state distribution in MDPs.
@@ B.3 Refined Coverage
B.4 Interpretation of 8
To interpret Assumption 8, let $x=\mathbf{u}(f_{h})/Z(f_{h})\in\Delta({\mathcal{S}})$, and assign a distribution over $x$ by sampling $f_{h}\sim\textrm{diag}(Z_{h}/S)$, then Assumption 8 becomes $x^{\top}\mathbb{E}_{Z_{h}/S}[xx^{\top}]^{-1}x\leq C_{\mathcal{F},U}S$ for any $x$ with non-zero probability. This is a common regularity assumption [Duan et al., 2020, Perdomo et al., 2023], requiring that no $x$ points to a direction in $\mathbb{R}^{S}$ alone without being joined by others from the distribution. While $1/\sigma_{\min}(\Sigma_{\mathcal{F},h})$ can bound this quantity, the other way around is not true, implying that Assumption 8 is weaker.
B.5 Intuition for $L_{\infty}$ Outcome Coverage
Here we provide more details about the $L_{\infty}$ outcome coverage in Section 4.2, which are omitted in the main text due to space limit.
Looseness of $L_{2}$ outcome coverage
We start by examining the looseness of $L_{2}$ outcome coverage (Assumption 7), which will motivate the $L_{\infty}$ version of outcome coverage. To develop intuitions, we first examine the boundedness of Eq.(4) (the construction that leads to $L_{2}$ outcome coverage) in the setting of Example 2. In this example, $\Sigma_{\mathcal{F},h}=\mathbf{I}$, and $\|V_{\mathcal{F}}\|_{\infty}\leq H$. According to the $L_{2}$ Hölder decomposition in Proposition 4, however,
The comparison clearly highlights that the looseness in $S$ comes from the fact that $L_{2}$ Hölder does not fully leverage the fact that $\|\mathbf{u}(f_{h})/Z(f_{h})\|_{1}=1$ and loosely relaxing it to $\|\mathbf{u}(f_{h})/Z(f_{h})\|_{2}\leq 1$. In contrast, our $L_{\infty}$ coverage assumption allows for a more natural $L_{1}/L_{\infty}$ Hölder to leverage the $L_{1}$ normalization of $\mathbf{u}(f_{h})/Z(f_{h})$.
$L_{\infty}$ Outcome Coverage
Similar to $L_{\infty}$ belief coverage, we could define $L_{\infty}$ outcome coverage simply as $\|(\Sigma_{\mathcal{F},h})^{-1}V_{{\mathcal{S}}}^{\pi_{e}}\|_{\infty}$. The small caveat and inelegance is that it does not recover $V_{\mathcal{F}}=R^{+}$ when $\pi_{b}=\pi_{e}$. To address this, we leverage the lesson from belief coverage (Section 5.2): to obtain $C_{\mathcal{H},\infty}=1$ when $\pi_{b}=\pi_{e}$, a key property is that $\Sigma_{\mathcal{H},h}$ (data covariance matrix) and $\mathbf{b}_{h}^{\pi_{b}}$ (the vector to be covered) are the covariance matrix and the mean vector w.r.t. the same distribution of belief vectors. In contrast, for the outcome coverage case, $V_{{\mathcal{S}}}^{\pi_{e}}$ depends on the reward function, but such information is missing in $\Sigma_{\mathcal{F},h}$, which prevents a perfect cancellation in the on-policy case.
Therefore, we can adjust the definition of $\Sigma_{\mathcal{F},h}$ to mimic the situation of belief coverage. The first step is to find the counterpart of belief state which is $L_{1}$-normalized. This obviously corresponds to $Z_{h}^{-1}M_{\mathcal{F},h}^{\top}$, whose rows sum up to $1$. Let $\bar{\mathbf{u}}(f_{h}):=\mathbf{u}(f_{h})/Z(f_{h})\in\Delta({\mathcal{S}}_{h})$, then $\Sigma_{\mathcal{F},h}=S\cdot\mathbb{E}_{Z_{h}/S}[\bar{\mathbf{u}}\bar{\mathbf{u}}^{\top}]$. Then, by incorporating reward information into $\Sigma_{\mathcal{F},h}$, we arrive at the solution in Eq.(5).
Finally, we examine the setting of Example 2 and show that $V_{\mathcal{F}}$ in Eq.(5) is similarly well-behaved as Eq.(4) in this scenario.
Example 9.
In the same setting as Example 2, i.e., $f_{h}$ always reveals $s_{h}$, we have $\Sigma_{\mathcal{F},h}^{R}=\textrm{diag}(V_{{\mathcal{S}}_{h}}^{\pi_{b}})$, and $\|(\Sigma_{\mathcal{F},h}^{R})^{-1}V_{{\mathcal{S}}}^{\pi_{e}}\|_{\infty}=\max_{s_{h}}V_{{\mathcal{S}}}^{\pi_{e}}(s_{h})/V_{{\mathcal{S}}}^{\pi_{b}}(s_{h})$.
This example also closely resembles Example 7 for belief coverage. While $(\Sigma_{\mathcal{F},h}^{R})^{-1}$ may behave poorly if $V_{{\mathcal{S}}}^{\pi_{b}}(s_{h})$ is small, an easy fix is to simply add a constant to the reward function and shift its range to e.g., [1, 2], which ensures a small $C_{\mathcal{F},\infty}$.
B.6 Extension to History-dependent Policies
The main text assumes memoryless $\pi_{b}$ and $\pi_{e}$. Similar to Uehara et al. [2022a], we can extend the approach to handle history-dependent policies (the changes to the proof are sketched below). Instead of rewriting the proof with these changes, we will provide a “black-box” reduction that handles the extension more elegantly and allow for more general results than Uehara et al. [2022a].
Memory-based Policies and Changes in the Proofs
Define memory $m_{h}$ as a function of $\tau_{h}$ (i.e., $m_{h}=m_{h}(\tau_{h})$), and a memory-based policy can depend on $(m_{h},o_{h})$. Uehara et al. [2022a] allows for $m_{h}=(o_{h-M:h-1},a_{h-M:h-1})$ for some fixed window $M$ in their analyses. If we let $m_{h}=(o_{1:h-1},a_{1:h-1})$, we recover fully general history-dependent policies.
If we direct modify the proofs to accommodate this generalization (as done in Uehara et al. [2022a]), the required changes are:
  • $V_{{\mathcal{S}}}^{\pi_{e}}$ and $V_{\mathcal{F}}$ need to additionally depend on $m_{h}$ to be well defined. $V\in\mathcal{V}$ also generally depend on $m_{h}$ since we need realizability $V_{\mathcal{F}}\in\mathcal{V}$.
  • $M_{\mathcal{F},h}$ and $M_{\mathcal{H},h}$ have rows indexed by $(s_{h},m_{h})$ instead of just $s_{h}$. That is, we replace belief state with posterior over $(s_{h},m_{h})$. Similarly, entries in $M_{\mathcal{F},h}$ are now future probabilities conditioned on $(s_{h},m_{h})$
However, if we consider latent state coverage, which is weaker than belief coverage that is really needed (Section E.3), we now require bounded $\frac{d^{\pi_{e}}(s_{h},m_{h})}{d^{\pi_{b}}(s_{h},m_{h})}$. This is generally exponential in the length of $m_{h}$ (e.g., if $M=h-1$, the ratio is exactly the cumulative importance weight), preventing us from handling $\pi_{b}$ and $\pi_{e}$ with long-range history dependencies.
Reduction to Memoryless Case
Instead of changing the proofs, we now describe an alternative approach that (1) produces results similar to Uehara et al. [2022a], and (2) handles more general recurrent policies that are finite-state-machines (FSMs), which subsume fully history-dependent policies (or policies that depend on a fixed-length window) as special cases.
Concretely, a policy with memory $m_{h}$ is said to be an FSM, if $m_{h+1}$ can be computed solely based on $m_{h},o_{h},a_{h}$, without using other information in $\tau_{h}$. $m_{h}=(o_{h-M:h-1},a_{h-M:h-1})$ satisfies this definition, as computing $m_{h+1}$ is simply dropping the oldest observation-action pair from $m_{h}$ and appending the newest one. [^6] When $h\leq M$ no dropping is needed. In fact, if we never drop, then $m_{h}=\tau_{h}$ is just the history. Another example is belief update, where $\mathbf{b}(\tau_{h+1})$ can be computed from $\mathbf{b}(\tau_{h}),o_{h},a_{h}$.
We assume that both $\pi_{b}$ and $\pi_{e}$ have the same memory; if they differ, we can simply concatenate their memories together. Then, handling memories in our analyses takes two steps:
  1. 1.
    We allow latent state transition to depend on $o_{h}$, that is, $s_{h+1}\sim\mathbb{T}(\cdot|s_{h},a_{h},o_{h})$. This model has been considered by Jiang et al. [2017] to unify POMDPs and low-rank MDPs. Our analyses hold as-is without any changes. To provide some intuition: the key property that enables the analyses of FDVF is that $s_{h}$ is a bottleneck that separates histories from futures, which enables $(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})=\langle\mathbf{b}(\tau_{h}),\mathcal{B}^{{\mathcal{S}}}_{h}V\rangle$ in Eq.(2). This property is intact with the additional dependence of $s_{h}$ on $o_{h-1}$.
  2. 2.
    We provide a blackbox reduction from the memory-based case to the memoryless case. Define a new POMDP that is equivalent to the original one, where the latent state is $\tilde{s}_{h}=(s_{h},m_{h})$. The observation is $\tilde{o}_{h}=(o_{h},m_{h})$, which can be emitted from $\tilde{s}_{h}$ since $\tilde{s}_{h}$ contains $m_{h}$. The latent state transition is $\tilde{s}_{h+1}=(s_{h+1},m_{h+1})$, where $s_{h+1}\sim\mathbb{T}(\cdot|s_{h},a_{h},o_{h})$, and $m_{h+1}$ is updated from $m_{h}$ (contained in $\tilde{s}_{h}$), $a_{h}$, and $o_{h}$; this is why we need $o_{h}$ to participate in latent transitions. Now it suffices to perform OPE in the new POMDP. Data from the original POMDP can be converted to that of the new POMDP with $\tilde{o}_{h}=(o_{h},m_{h})$. $\pi_{b}$ and $\pi_{e}$ only need to depend on $\tilde{o}_{h}$ and become memoryless.
Given the reduction, if we design function classes that operate on the histories and futures of the new POMDP, the guarantees in the main text immediately hold. The final step is to translate the objects (e.g., futures and histories) and guarantees in the new POMDP back to the original POMDP for interpretability. Note that the history in the new POMDP is $\tilde{\tau}_{h}=(\tilde{o}_{1:h-1},a_{1:h-1})$, which contains the same information as $\tau_{h}$, so functions in $\Xi$ and $\mathcal{W}$ can still operate on $\tau_{h}$. Similarly, $V\in\mathcal{V}$ now takes $(m_{h},f_{h})$ as input, which is consistent with Uehara et al. [2022a].
When translating the assumptions back to the original POMDP, we can see that now the results can be well-behaved when $m_{h}$ is “simple”. For example, if $m_{h}$ takes values from a constant-sized space, $d^{\pi_{e}}(s_{h},m_{h})/d^{\pi_{b}}(s_{h},m_{h})$ (which lower-bounds belief coverage as discussed above) may not blow up exponentially, even though $m_{h}$ can hold information that is arbitrarily old (e.g., it remembers one bit of information from $h=1$). This is a scenario that cannot be handled by the formulation of Uehara et al. [2022a].
[…]
B.7 Recovering MDP Algorithms and Analyses
One somewhat undesirable property of our algorithms and analyses is that they do not subsume MDP algorithms/analyses as a special case. MDPs can be viewed as POMDPs with identity emission, i.e., $o_{h}=s_{h}$. In this case, algorithms considered in this paper are analogous to their MDP counterparts Uehara et al. [2020], except that the MDP algorithms require that all functions $V$, $w$, $\xi$ to operate on the current state $s_{h}$. In contrast, our FDVF operates on $f_{h}=(o_{h},a_{h},\ldots,o_{H},a_{H})$ which includes $o_{h}$($=s_{h}$). To recover the MDP algorithm as a special case, we can choose $\mathcal{V}$ such that every $V\in\mathcal{V}$ only depends on $f_{h}$ through $o_{h}$. However, $w$ and $\xi$ operate on $\tau_{h}=(o_{1},a_{1},\ldots,o_{h-1},a_{h-1})$ which does not contain $o_{h}$. This makes subsuming the MDP case difficult.
Here we describe briefly how to overcome this issue by slightly modifying our analysis; the changes are somewhat similar to Appendix B.6. Instead of letting $\tau_{h}=(o_{1},a_{1},\ldots,o_{h-1},a_{h-1})$, we can define an alternative notion of history $\tilde{\tau}_{h}=(o_{1},a_{1},\ldots,o_{h})$ and replace $\tau_{h}$ in the main text with $\tilde{\tau}_{h}$. Algorithmically, we can immediately recover MDP algorithms by also restricting functions in $\mathcal{W}$ and $\mathcal{X}$ to only operate on $o_{h}$. However, our analyses (which apply to general POMDPs) need to change accordingly, as replacing $\tau_{h}$ with $\tilde{\tau}_{h}$ will break the key properties, such as $\mathcal{B}^{\mathcal{H}}V$ being linear in $\mathcal{B}^{{\mathcal{S}}}V$. The problem is that in the definitions of $V_{{\mathcal{S}}}^{\pi_{e}}$ and $\mathcal{B}^{{\mathcal{S}}}V$ we want to marginalize out the randomness of $o_{h}$ given $s_{h}$, which is in conflict with conditioning on $\tilde{\tau}_{h}$ that includes all the information of $o_{h}$. To address this, we simply replace $s_{h}$ with $\tilde{s}_{h}:=(s_{h},o_{h})$ in the definition of $V_{{\mathcal{S}}}^{\pi_{e}}$, $\mathcal{B}^{{\mathcal{S}}}V$, $M_{\mathcal{H},h}$, and $M_{\mathcal{F},h}$. That is, $V_{{\mathcal{S}}}^{\pi_{e}}$ is now a function of $\tilde{s}_{h}$ and also depends on $o_{h}$, $\mathbf{b}(\tau_{h})$ is the posterior distribution over $\tilde{s}_{h}$, and the entries of the outcome matrix $M_{\mathcal{F},h}$ is $\Pr_{\pi_{b}}(f_{h}|s_{h},o_{h})$. This retains the key property that $\mathcal{B}^{\mathcal{H}}V$ is linear in $\mathcal{B}^{{\mathcal{S}}}V$ with $\tilde{\mathbf{b}}(\tau_{h})$ as the coefficient.
When we specialize the guarantees of this modified analysis to the MDP setting, outcome coverage is always satisfied and we can always use the MDP’s standard value function as $V_{\mathcal{F}}$. For belief coverage, note that $(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})=(\mathcal{B}^{{\mathcal{S}}}V)(s_{h},o_{h})$, which is simply the standard definition of Bellman error in MDPs (since $s_{h}=o_{h}$), which we denote as $(\mathcal{B}V)(s_{h})$. Plugging this into the refined coverage discussed in Section B.3, we recover the standard definition of coverage in the MDP setting, namely $\sup_{V\in\mathcal{V}}\frac{|\mathbb{E}_{\pi_{e}}[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})]|}{\sqrt{\mathbb{E}_{\pi_{b}}[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})^{2}]}}$.
B.8 Incorporating Different Latent-State Priors in $V_{\mathcal{F}}$
In Section 4 we showed that $\Sigma_{\mathcal{F},h}$, which is crucial to the construction in Eq.(4), can be viewed as the confusion matrix of making posterior predictions of $s_{h}$ from $f_{h}$, using a uniform prior. The uniform prior corresponds to the all-one vector $\mathbf{1}_{{\mathcal{S}}}$ in the definition of $Z_{h}:=\operatorname{diag}(\mathbf{1}_{{\mathcal{S}}}^{\top}M_{\mathcal{F},h})$, which naturally leads to the question of whether we can incorporate a different and perhaps more informative prior.
Let $\mathbf{p}_{h}\in\Delta({\mathcal{S}}_{h})$ be the prior we would like to use instead of the uniform prior. An immediate idea is to define $Z_{h}^{\mathbf{p}_{h}}=\operatorname{diag}(\mathbf{p}_{h}^{\top}M_{\mathcal{F},h})$ (possibly up to a $S$ scaling factor, as the uniform prior is $\mathbf{p}_{\textrm{unif}}=[1/S,\cdots,1/S]^{\top}$ and $\mathbf{1}_{{\mathcal{S}}}=S\mathbf{p}_{\textrm{unif}}$) and directly plug it into the construction in Eq.(4). However, this breaks some of the key properties of the current construction of $\Sigma_{\mathcal{F},h}$, such as $L_{1}$ normalization of rows of $Z_{h}^{-1}M_{\mathcal{F},h}^{\top}$.
To resolve this, the key is to realize that the rows of $Z_{h}^{-1}M_{\mathcal{F},h}^{\top}$ are $L_{1}$ normalized because they can be viewed as posteriors over ${\mathcal{S}}_{h}$. In comparison, if we examine the $(f_{h},s_{h})$-th entry of $(Z_{h}^{\mathbf{p}_{h}})^{-1}M_{\mathcal{F},h}^{\top}$, it is
$\frac{\Pr_{\pi_{b}}[f_{h}|s_{h}]}{\sum_{s^{\prime}\in{\mathcal{S}}_{h}}\Pr_{\pi_{b}}[f_{h}|s^{\prime}]\mathbf{p}_{h}(s^{\prime})},$
so clearly it is missing a $\mathbf{p}_{h}(s_{h})$ term on the numerator to be a proper posterior. Inspired by this observation, we can see how to fix the construction now: recall that $V_{\mathcal{F}}$ needs to satisfy $M_{\mathcal{F},h}V_{\mathcal{F},h}=V_{{\mathcal{S}},h}^{\pi_{e}}$. If $\mathbf{p}_{h}>0$, then this is equivalent to
$(\operatorname{diag}(\mathbf{p}_{h})M_{\mathcal{F},h})V_{\mathcal{F},h}=\operatorname{diag}(\mathbf{p}_{h})V_{{\mathcal{S}},h}^{\pi_{e}}.$
Then, the minimum $Z_{h}^{\mathbf{p}_{h}}$-weighted solution of this equation will provide the desired construction that preserves the $L_{1}$ normalization properties.
Which prior $\mathbf{p}_{h}$ to use?
For the initial time step $h=1$, the initial latent-state distribution $d_{1}$ is the most natural candidate for the prior. If $d_{1}(s_{1})=0$ for some $s_{1}\in{\mathcal{S}}_{1}$, incorporating $d_{1}$ as the prior will essentially treat $s_{1}$ as non-existent and ignore the outcome coverage of $\pi_{b}$ over $\pi_{e}$ from $s_{1}$, which is reasonable because $s_{1}$ will not be activated by neither $\pi_{e}$ and $\pi_{b}$. For $h>1$, the answer is less clear, and natural candidates include $\mathbf{p}_{h}=d^{\pi_{b}}_{h}$ and $d^{\pi_{e}}_{h}$, or perhaps their probability mixture. For the example scenarios examined in the main text, these choices (or even a uniform prior) do not make significant differences, and we leave the investigation of which prior is the best to future work.
Appendix C Proofs for Section 3
C.1 Proof of Lemma 1